Transversity/pre-beam checklist

From Hall A Wiki
Jump to: navigation, search

This is Transversity Experiment Pre-Beam Checklist.

Xiaodong Jiang 07/21/2008, Updated 07/23/08 after the collaboration meeting.

(Please don't erase items. Add comments to the items when they are done. Add new items at the end of each section. Please highlight the changes you made.)

1. Beamline,beam delivery. Compton, Lumi-detectors and beam-dump.

   1.1 Transversity beam delivery specifications, communication to MCC 
       (Xiaodong Jiang + J.-P. Chen with MCC).

   1.2 Beam ramp-up procedures on polarized He3 target.
       (J.-P. Chen with MCC).

   1.3 Confirm final beam energy for delivery (one- two- and five-pass),
       update runplan on kinematics settings (Xiaodong Jiang). 

   1.4 Hall A harp check out, confirm harp-scan procedures (Arun Saha).

   1.5 Hall A BPM check out. BPM calibration need and procedures. 
       Calibration constants and online software checkout for spot++ (XXX.XXX)

   1.6 Decide raster pattern: SQUARE (Done! 7/21/08 J.-P. Chen),
       specify beam size on polarized He3 cell (2mm by 2mm to 4mm by 4mm),
       raster hardware check (Hall A staff), software control and decoding in place (Alex).

   1.7 Compton detector checklist (Sirish Nanda and CMU group).
       Beam tuning procedures for Compton, specify the acceptable beam tune (Sirish with MCC).
       Compton commissioning run plan.
       Setup for Compton running during production, local experts for trouble shooting.

   1.8 Moller preparation for one- and five-pass beam (Eugene Chudakov).
       Production runs with Moller quards ON for five-pass (Yes, that's the default, Eugene). 
       Consistecy checks with Hall B and Hall C Moller measurements.

       Determine absolute beam helicity signs in Hall A, for one- and 
       five-pass beam. Runs with a large beam charge asymmetry (Yes, we need to do that. Xiaodong Jiang). 

       Beam half-wave plate IN/OUT changes during data taking, need to communicate with other halls.

       Wiene filter angle and beam half-wave plate status in EPICS data stream (Yes. Eugene).
   1.9 Set trip limits for target and beam-dump ion chambers.

   1.10 Lumi detector checkout, remove neutral density filters (Dustin McNulty). 
       HAPPEX DAQ, beam charge feedback checkout.
   1.11 Lumi readout by transversity DAQ. Synchronization and consistency 
        check between Lumi, HRS scalers and BigBite scalers 
       (Bob Michaels + Joe + Kalyan). 

   1.12 An iron (?) cylinder mounted around the first sections of downstream beam pipe (Doug).

   1.13 Procedures for testing BigBite's steering of beam at one-pass (Doug). 

   1.14 Beamline survey, relations to target (Hall A staff in charge ?).

   1.15 Beam charge monitors checkout, fix down*3 problem (Arun Saha).
        Confirm calibration constants in online software (Arun Saha).

        Need a beam charge calibration in Hall A.
        Prepare a procedure and checkout for hardware.

2. Targets.

  2.1 Target magnetic field measurement (Yi Qiang)

  2.2 Target holding field direction measurements with compass.

  2.3 Target cell alignment and survey. 

  2.4 Target position decoding, calibration of target movement.

  2.5 NMR and EPR working.

  2.6 Target cooling jet working.

  2.7 Target slow control variables in EPICS data stream.
      Target temperture sensors working and calibrated.

  2.8 Target spin-flip tested, integrated with DAQ.
      Nominal expected target polarization with 20 minutes target spin flip (?).
      Keep flipping target spin regardless of beam and DAQ status (Yes. J.-P. Chen).
      Roughly balence the integrated beam charge for spin-up vs spin-down runs(?).

      A detailed commisioning run plan and checklist for spin-flip test with beam.
      (Yi Qiang and Jin Huang. Trans. in-plane polarization, one-pass long. polarized  beam 
       double-spin asymmetry, at Delta-production kinematics). 

  2.9 Reference cell leak check (Todd).
      Reference cell control tested for He3, H2, N2 gas, 
      pressure transducers cross-calibrated (Todd)
  2.10 Target safty and laser safty documents in place (J.-P. Chen).

  2.11 Procedures of broken cell recovery, agreed with RADCON (J.-P. Chen).

  2.12 A checklist for laser optics.

  2.13 Rotate target to "Long+" and "Long-" for elastic He3 runs at one-pass (Yes, Xiaodong Jiang). 

3. Let HRS.

  3.1 Left HRS free to move between 16 degree and ~60 degree.
      Pins on the floor to stop it from moving beyond 15.5 (?) degree.
  3.2 Magnet control.  Q3 control fixed. Quards cycling procedures.

  3.3 Procedure of polarity change, minimize time taken (Ed Folts).

  3.4 Sieve slit insert ready. Survey positions. Reproducebility of location (Doug).

  3.5 Front vaccum window cover ON before beam comes (Ed Folts).
      Front vaccum window cover removed right before taking data (Ed Folts, checklist).   

  3.6 Detector cosmic ray runs (Vince+Alex+INFN).
      All detector channels checked.
      A final layout of HRS detectors, specify their postions (Vince with Al).

  3.7 Timing offsets of S1 and S2M determined.
      Cosmic ray beta spectrum make sense. (Vince).

  3.8 Verify the initial optics data base for extended target (Vince).

  3.9 Initial time-of-flight corrections in place (Vince).

  3.10 RICH commissioning software ready.  Prepared to murge into Hall A analyzer.
       (INFN: Guido Urciuoli)

  3.11 A1 detector PMT HV gain matched. Software gain-match tool ready.
       Check dry gas purging (Vince+YouCai).

  3.12 Short gas Cherenkov PMT HV gain-matched. Software gain-match tool ready

  3.13 Lead glass PMT HV gain matched, software gain-match tool ready (Vince).

  3.14 A detailed commisioning run plan in place for each detector.
       Specify the conditions to be satisfied before moving into production (Vince).

4. BigBite.

   4.1 A final detector configuration plot, 
       with each detector's position laid out (Doug with Al).

   4.2 Verify BigBite's final position in the hall. 
       Survey of BigBite and the detector positions (Doug with JLab survey group).

   4.3 BigBite dipole magnet control, work to 710 A.
       BigBite magnetic probe in EPICS data stream (Doug).
       With a Hall probe, measure the maximum residue magnetic field at the detector's PMT locations
       when teh magnetis on, and the iron shielding house is in place.  Verify that the field is small.  

   4.4 BigBite dipole magnet insert pieces in place, both on the bottom 
       and at the top (Doug with Ed Folts).

   4.5 Front sieve slit mounted, moving mechenisim and its reproducebility 
       checked. Sieve slit position surveyed.

   4.6 Detector package cosmic ray checked.
       Clearly show three-chamber tracking correlated with lead-glass hits.

   4.7 Wire chamber HV control. 
       Threshold voltage control works from the counting house. 
       Voltages and other variables into the EPICS data stream.

   4.8 Gas checkout procedures for wire chambers, gas Cherenkov.
       Why not flow gas directly from outside Hall A ? (Jack)
       Connect dry N2 gas and flush PMT housing. 
   4.9 Detecors' relative positions surveyed (Doug). 
       Initial position offsets put into analyser (Jin Huang).

   4.10 Intitial optics data base set up for 1.5 meter drift (Xin Qian).
       Software tools ready for analysing optics data (Xin Qian and Jin Huang).

   4.11 Scintillator time offsets determined. 

       Initial flight path corrections for time-of-flight 
       in place for electrons in BigBite, be able to identify beam bunches with relative timing 
       through beam RF (Vince).

   4.12 Add shielding blocks to protect front-end electronics from radiation damage 
        (on both upstream and downstream).
        Add shielding blocks to teh BigBite DAQ weldment on the upstream side.      
5. DAQ, time-of-flight, scalers, slow control and  online scripts.

   5.1 Time-of-flight logic pulser checked.
       Verify calculation of intrinsic delays for H(e,e'p) 
       one-pass and two-pass beam 

       Verify intrinsic delays for production kinematics.

   5.2 Beam RF timing signal readout (was in data for lowQ2-ep, Vince).

   5.3 High resolution TDC for BigBite scintillators (LeCroy 1875 TDC obtained 
       through Doug 07/22/08, 50 ps resolution). 
       Checkout a spare 1875 TDC module.

       A list of actions/tests for shortening the time-of-flight window
       at the beginning of production. 

   5.4 Cosmic gain-match checked for BigBite pre-shower and shower.

   5.5 Remote control of trigger threshold.
       Two-level trigger threshold settings, a low and a high (?). 

   5.6 Scalers working (Kalyan + Bob Michaels). 
       Scaler display working in the counting house. 

   5.7 Online dead time calculation make sense from the scalers.
       Minimize DAQ readout time, reduce deadtime (Bob Michaels + Kalyan).

   5.7 EPICS variables into the data stream.
       A list of online EPICS charts to watch by the shift workers.

   5.8 Modify: Beginning-of-run script and End-of-run script.
       Automatic entry to Halog.

   5.9 Raw data file size limit (8 GB now ? need final confirmation, Ole), no file splitting.
       Scalers won't over flow (?) Take 100 kHz clock, would the analyzer handle a scaler wrap-up (?)

       File name convention: e06010_xxxx.dat.

       Online data disk spaces (Ole). 
       Scripts for copying raw data file to silo and verify that no recorded data get lost (Bob Michaels).
       A backup copy for raw data (Bob Michaels).

   5.10 A complete detector-map file (Jin Huang).

6. Documentations and on-line instructions.

   6.1 Run sheet and updates (Xiaodong Jiang). 

   6.2 Shift checklist, changes and updates (Xiaodong Jiang).

   6.3 Shift worker's how-to webpage, set up for transversity and update often.

   6.4 Shift worker training, Hall A walk-through.
       Target operators training and certification (J.-P. Chen).

   6.5 Shift worker rules, 
       and shift worker's food in the counting house.

   6.6 Experiment documentations, RSAD (done), ESAD, COO etc (J.-P. Chen).
       Folders in the counting house.

   6.7 A final Hall A floor layout for Transversity (Xiaodong Jiang with Al).

   6.8 Cable routes and floor layout (Brad).

7. Transversity start-up conditions on Oct.7th, 2008.

  7.1 One-pass beam to Hall A, E0=1.234 polarized.

  7.2 Left HRS at 16 degree, negative mode, quards cycled.
      P_HRS=1.174 GeV/c for C12 quasi-elastic kinematics.
      Front sieve slit insert:OUT.

  7.3 BigBite at 30 degree, beam right. 1.5 meter drift distance 
      from center of target to the front surface of BigBite magnet.
      Front sieve slit: OUT.

      Magnet set to NEGATIVE mode, with 710 A current.

  7.4 Target position: EMPTY.

      Ready to move to: 
       Carbon hole,
       Carbon multi-foil,
       Reference cell.

8. Services and responsibilities

  8.1 Safty and experiment documents.

      RSAD (Done)
      ESAD, COO, etc (J.P. Chen, deadline: 10 days before Oct. 7th, 2008).

  8.2 Wiki webpage: policing, clearing-up and backing-up (Yi Qiang).

  8.3 Traneversity collaboration web page, related physics papers and 
      collaboration meeting records (Kalyan). 

  8.4 Shift sign-up script setup (Alex).
      Daily shift changes and updates (Xiaodong Jiang + Jin Huang).

      Experts' shifts at commisioning, and production (Xiaodong Jiang).

  8.5 Run plan and updates (Xiaodong Jiang).

  8.6 Run coordinators and backup person update (Xiaodong Jiang).

  8.7 A list of contact for technical experts (Jin Huang).

      A list of collaborators and their institutions, as shifts being
      singed-up (Jin Huang and Xiaodong Jiang).

  8.8 A list of good runs with comments, daily updates (XXX.XXX)
      Keep track of beam-charge, number-of-good-events.

      Decide when to change HRS polarity, target spin orientations.

  8.9 A "lay level" description of transversity experiment for JLab public webpage (Jen-Chieh Peng) 

9. Online replay.

   9.1 Jin Huang is in charge with Vince's suppervision. 
   9.2 Mantain the official version of transversity analyzer.
       Software setup, update and backup files. 
       Online instructions. 
       Standard spectrum for online shift workers.
       Standard online data quality cuts and keep track the number of good events (Jin Huang). 

   9.3 Semi-Inclusive DIS physics class with kinematic variables (Xin Qian).
       Online display of physics variables (Xin Qian).
       Online calculation of inclusive channel double-spin asymmeries for monitoring (Xin Qian)

10. Spare parts and emergency procedures.

  10.1 Two spare LeCroy HV crate tested and working.

  10.2 Spare wire chamber HV modules tested.

  10.3 Two spare FASTBUS crates and power supplies tested (Doug with JLab electronic group). 

       Network cards.

  10.4 Spare threshold power supplies for BigBite wire chambers (Xin Qian with Jack).
       Spare PreAmp cards, level translators, long and short ribbon cables (Xin Qian). 
       Spare LeCroy 1877 TDCs tested (Xin Qian).

  10.5 A list of target spare parts (Yi Qiang).

  10.6 Spare lasers and spare optics fibers.

  10.7 Working target cells. Referece cells.

       Procedures of replacing cells and re-alignment (J.-P. Chen).

  10.8 Procedures of recovery after power failures, hurricans.

       Save initial setup files.
       Restart GUI.
       Reboot IOCs, control computers, target computer. Restart spin-flip.
       Emergency contact list.

  10.10 At what point we should consider to give up.

        Too bad we can't have production runs because:

        Target can not be polarized to 20% polarization.
        Spin-flip doesn't work.

        No coincidence T5 triggers or can't see coin-TOF peak.
        (Double check coincidence  rates with Hall C semi-inclusive data at x=0.3)
        (Xiaodong Jiang and Kalyan Allada, specify the range of acceptable rates).
        DAQ is too slow, can't take 100 Hz coincidence trigger rate.
        BigBite trigger doesn't work.
        Wire chambers can not operate, too much background or can't find any clear.

        Chambers are too noisy, raising threshold on PreAmp cards doesn't work.
        Chambers have more than 6 (?) broken planes in total.

        FASTBUS is not working, No spare parts.       

        Unknown DAQ problems. Error flags can not be understood. 

        Beam trips too often at five-pass (more than 10 times an hour).

  10.11 After the 10-days commisioning period, extensive repair work on the 
        following items have to wait till the next scheduled down time:

        i. BigBite gas Cherenkov.
       ii. Individual pre-shower and shower blocks.
      iii. Less than three wire planes in BigBite chambers
       iv. HRS RICH detector.
        v. Replacement of PMT in HRS gas Cherenkov or A1.   
       vi. Less than four failed Lumi-detctors.
      vii. Compton related items.

  A Timeline for Transversity Preparation Work

  1. Oct. 7th: All sub-systems ready to take beam.

  2. Oct. 1st: Integrated test starts for polarized target and DAQ. 
  3. Sept. 15th: Cosmic ray test for left-HRS including RICH.
     Coincidence trigger tested with a pulser.

     BigBite detectors in their final place, all cables connected.
     Lumi detectors readout in DAQ.
     Scalers working.

  4. Sept. 1st: BigBite wire chamber readout test done, all channels working.
                BigBite gas Cherenkov installed into the detector package. 
                HRS detector read out done without RICH. 
  5. Aug. 15th: BigBite gas Cherenkov's problem understood, solutions worked out.
                Missing mirror installed.

                BigBite wire chambers readout test started.
                Polarized target field measurement finished.

  6. Aug. 1st: BigBite wire chambers installed into detector package, positions 

               Target coil installation finished.

               All target design finalized. Parts on order.

  7. July 21st: Collaboration meeting.

                Identify all missing items. Clearify responsibilities.
                A list of designs need to be finished.

                A final preparation schedule for all sub-systems to follow. 
                Update the installation schedule.